OpenStudy (anonymous):

If y=6 when x=9 , find x when y=12

OpenStudy (anonymous):

How did u get 8

hero (hero):

Do you know if you're dealing with inverse variation or direct variation?

OpenStudy (anonymous):

Direct variation

hero (hero):

The formula for direct variation is y = kx In the first case, when x = 9 and y = 6, you have 6 = 9k 6/9 = k 2/3 = k In the second case, they give you y = 12, and ask you to find x. Now k is constant so: 12 = (2/3)x Thus x = 12(3/2) x = 18

OpenStudy (anonymous):

Thanks